# DeepExtractor: Time-domain reconstruction of signals and glitches in gravitational wave data with deep learning

### v1

@@ V Conclusions and Future Work
The key innovation lies in the training strategy. Rather than directly mapping input data to clean waveforms, the model is trained to isolate the noise component from input data comprising background noise and injected signals or glitches. By subtracting the predicted noise, DeepExtractor reconstructs the signal or glitch. This approach outperforms direct mapping in generalization and underscores the importance of accurately modeling the detector’s power spectral density (PSD). The framework is highly flexible, capable of reconstructing any power excess above the inherent detector noise. This flexibility stems from training on a diverse dataset of linear combinations of waveforms from five distinct analytical classes, enabling effective interpolation to reconstruct unseen waveforms.
Three experiments validated DeepExtractor’s efficacy. The first experiment involves simulated Gaussian noise with injected glitches, measuring time-domain match performance across several deep learning models. The results demonstrate that our U-Net architecture significantly outperforms autoencoders without skip connections, even when the overall architectures are otherwise identical. Additionally, U-Nets operating on magnitude and phase spectrograms are shown to deliver superior performance compared to a 1D U-Net applied directly to the corresponding time-series data.
@@
Finally, we evaluated DeepExtractor’s capability to reconstruct real glitch events from LIGO’s third observing run. This represents the first comprehensive application of deep learning to glitch reconstruction. By utilizing Q-scan and time-domain visualizations, we ensured effective glitch removal while preserving the underlying noise characteristics. The study highlighted the importance of integrating PSD knowledge through transfer learning, allowing the model to adapt effectively to the complexities of real detector data.
For signal reconstruction, a key next step is to compare DeepExtractor’s performance against established search and parameter estimation pipelines. Currently, the model does not handle overlapping signals and glitches, but this limitation could be addressed by extending the training framework. Instead of isolating only the noise component, the model could be trained to separate noise and signals from data containing noise, signals, and glitches. Simulated signal components could be generated using waveform models commonly employed in parameter estimation. Incorporating data from multiple detector channels could further enhance performance, particularly in scenarios with overlapping signals and glitches.
@@ VI Acknowledgements
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by the National Science Foundation Grants No. PHY-0757058 and No. PHY-0823459.
@@ VII.1 Simulated training glitches
sine-gaussian Generates a sine wave modulated by a Gaussian envelope. Frequency and tau are randomly chosen. freq_min, freq_max, duration, sample_rate
@@ VII.2 STFT Parameters
VII.3 Reconstruction of unseen test classes
FIG. 11 shows examples of full 222\,2s extracted signals and zoomed-in views (middle 0.50.50.5\,0.5s) for classes that were unseen during training.
(a) Gengli H1
(b) cDVGAN Tomte
(c) cDVGAN BBH
(d) cDVGAN Simplex
@@ VII.3 Reconstruction of unseen test classes
VII.4 Results of Shapiro Tests on Gravity Spy data
Table 4 shows the mean and standard deviation of Shapiro-Wilk test p-values before and after mitigation via DeepExtractor over 100 random samples per class. Higher p-values after analyzing with DeepExtractor indicate that the output looks more Gaussian.
Category Before After
Extremely_Loud 0.00 $\pm$ 0.00 0.39 $\pm$ 0.33
Scattered_Light 0.49 $\pm$ 0.30 0.52 $\pm$ 0.29
Scratchy 0.36 $\pm$ 0.31 0.46 $\pm$ 0.27
Power_Line 0.43 $\pm$ 0.33 0.50 $\pm$ 0.32
Koi_Fish 0.00 $\pm$ 0.00 0.46 $\pm$ 0.30
Whistle 0.26 $\pm$ 0.34 0.50 $\pm$ 0.31
Low_Frequency_Burst 0.46 $\pm$ 0.30 0.47 $\pm$ 0.29
Tomte 0.16 $\pm$ 0.23 0.45 $\pm$ 0.28
Repeating_Blips 0.04 $\pm$ 0.17 0.44 $\pm$ 0.32
Fast_Scattering 0.49 $\pm$ 0.31 0.50 $\pm$ 0.30
Blip 0.03 $\pm$ 0.12 0.49 $\pm$ 0.28
Blip_Low_Frequency 0.37 $\pm$ 0.31 0.42 $\pm$ 0.31
Violin_Mode 0.13 $\pm$ 0.23 0.39 $\pm$ 0.27
Helix 0.02 $\pm$ 0.10 0.41 $\pm$ 0.31
Low_Frequency_Lines 0.47 $\pm$ 0.31 0.47 $\pm$ 0.31
Wandering_Line 0.17 $\pm$ 0.24 0.35 $\pm$ 0.33
No_Glitch 0.38 $\pm$ 0.30 0.41 $\pm$ 0.30
Air_Compressor 0.52 $\pm$ 0.28 0.53 $\pm$ 0.28
Chirp 0.24 $\pm$ 0.31 0.40 $\pm$ 0.28
Light_Modulation 0.00 $\pm$ 0.02 0.54 $\pm$ 0.28
Paired_Doves 0.29 $\pm$ 0.32 0.47 $\pm$ 0.30
1400Ripples 0.37 $\pm$ 0.35 0.49 $\pm$ 0.31
1080Lines 0.52 $\pm$ 0.32 0.57 $\pm$ 0.31
Table 4: Mean Shapiro-Wilk test p-values before and after glitch mitigation via DeepExtractor.
VII.5 Extracted Gravity Spy Glitches
Figures 12, 13, 14, 15 and 16 show DeepExtractor reconstructions for glitch classes from the Gravity Spy dataset.
(a) Air Compressor (03a, L1)
(b) Blip (03b, H1)
(c) Blip Low Freq. (03b, L1)
@@ VII.5 Extracted Gravity Spy Glitches
(a) Light Modulation (03a, L1)
(b) Low Frequency Burst (03b, H1)
(c) Paired Doves (03b, H1)
Figure 14:
(a) Power Line (03a, L1)
(b) Repeating Blips (03a, H1)
(c) Scattered Light (03a, H1)
Figure 15:
(a) Scratchy (03b, L1)
(b) Wandering Line (03b, H1)
(c) Whistle (03a, H1)
Figure 16:

### v3

@@ V Conclusions and Future Work
The key innovation lies in the training strategy. Rather than directly mapping input data to clean training waveforms, the model is trained to isolate the noise component from input data comprising background noise and injected signals or glitches. By subtracting the predicted noise, DeepExtractor reconstructs the signal or glitch. This approach outperforms direct mapping in generalization. The framework is highly flexible, capable of reconstructing any power excess above the inherent detector noise. This flexibility stems from training on a diverse dataset of linear combinations of waveforms from five distinct analytical classes (proxy glitches), enabling effective interpolation to reconstruct unseen waveforms.
Four experiments validated DeepExtractor’s efficacy. The first experiment involves simulated Gaussian noise with injected glitches, measuring time-domain mismatch performance across several deep learning models. The results demonstrate that our U-Net architecture significantly outperforms autoencoders without skip connections, even when the overall architectures are otherwise identical. Additionally, U-Nets operating on magnitude and phase spectrograms are shown to deliver superior performance compared to a 1D U-Net applied directly to the corresponding time-series data.
@@
We evaluated DeepExtractor’s capability to reconstruct real glitch events from LIGO’s third observing run. This represents the first comprehensive application of deep learning to glitch reconstruction. By utilizing Q-scan and time-domain visualizations, we ensured effective glitch removal while preserving the underlying noise characteristics. The study highlighted the importance of integrating PSD knowledge through transfer learning, allowing the model to adapt effectively to the complexities of real detector data.
Finally, we applied DeepExtractor to three real gravitational-wave events from LIGO’s third observing run, demonstrating its potential as a model-agnostic tool for signal reconstruction. The resulting reconstructions showed mismatches under 10% against the most of the parameter estimation templates, even without being explicitly trained on GW signals. Some of the observed mismatch is likely attributable to differences between our preprocessing pipeline and that used in parameter estimation. Aligning the preprocessing steps more closely with standard parameter estimation workflows, and incorporating signal models during training, may further enhance the reconstruction performance.
Several avenues remain open for future research. For example, modifying the training classes or tuning associated hyperparameters—such as the duration range of $[0.125,2]\,$s—could help assess whether these adjustments yield improved reconstruction performance. While BayesWave remains the most widely used tool for glitch mitigation in the gravitational-wave community, future work could compare DeepExtractor with alternative approaches, such as gwsubtract, specifically in cases where glitches exhibit linear correlations with auxiliary channels. Comparisons with other deep learning–based signal reconstruction methods, such as AWaRe, are also of interest. However, to be applicable for arbitrary glitch reconstruction, such models would need to be adapted to incorporate DeepExtractor’s training scheme—or a comparable approach capable of recovering excess power directly from strain data.
For signal reconstruction, a key next step is to compare DeepExtractor’s performance against established search and parameter estimation pipelines. While the model can reconstruct both a signal and glitch when they overlap in the data, it does not yet support their separation. This limitation could be addressed by extending the training framework to separately reconstruct overlapping components. Rather than isolating the noise only, the model could be trained to also output glitch and signal estimates, using simulated signals from standard waveform models as supervised targets. Furthermore, incorporating data from multiple detectors would allow DeepExtractor to model coherent signal power across the network, potentially improving reconstruction accuracy in cases where glitches and signals coincide.
@@ VI Acknowledgements
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by the National Science Foundation Grants No. PHY-0757058 and No. PHY-0823459. Finally, we thank the anonymous referees for their useful suggestions in improving the manuscript.
@@ VII.1 Simulated training glitches
sine-gaussian Generates a sine wave modulated by a Gaussian envelope. freq_min, freq_max, duration, sample_rate
@@ VII.2 STFT Parameters
VII.3 Test samples generated by gengli and cDVGAN
FIG. 12 and 13 show samples generated by gengli and cDVGAN respectively, used to evaluate the generalization capability of DeepExtractor. By interpolating across the class space, cDVGAN enables the creation of a diverse dataset that captures key features characteristic of both gravitational-wave signals and glitches. It is observed that by sampling the class vector, we can achieve a wide variety of morphologies that mix the features of the original training classes.
Figure 12: Two example Blip glitches from the Hanford (top row) and Livingston (bottom row) detectors, generated using gengli.
Figure 13: Examples of in-class (vertex) and class-mixed samples (simplex, uniform) generated by cDVGAN. The class vector used to generate each sample is shown at the top of each plot.
VII.4 Extracted Gravity Spy Glitches
FIG. 14, 15 and 16 show DeepExtractor reconstructions for glitch classes from the Gravity Spy dataset. Each figure shows reconstructions for three glitch classes which are named above each column. The top row shows a Q-scan of the input to the network. The middle row shows a Q-scan after removing DeepExtractor’s reconstruction. The bottom plots show the time series of the input and the reconstruction. The maximum color limit in the Q-scans is set to 25, similarly to the Q-scans shown in FIG. 9.
@@ VII.4 Extracted Gravity Spy Glitches
VII.5 Reconstructing O3 GWs with glitches
FIG. 17 shows DeepExtractor reconstructions for the same events shown in Section IV.4 but with blip glitches injected close to the merger using gengli. It is observed that when glitches are coincident with GWs, that DeepExtractor will reconstruct both sources of excess power without distinguishing between them.
(a) GW190521_074359
(b) GW200129_065458
(c) GW200224_222234
Figure 17: DeepExtractor reconstructions of gravitational-wave signals from the same three O3 events shown in Section IV.4, now with blip glitches injected near the merger in each detector. Glitches are injected with SNRs comparable to the matched-filter SNR of the corresponding maximum likelihood waveform. Each panel displays the model reconstruction overlaid with the aligned template and injected glitch. The residual plot compares the reconstruction to the linear combination of the signal and glitch.